I'm having problems trying to solve this equation. I think it's the brackets that are confusing me.
25 - [2 + 5y - 3(y + 2)]= -3(2y - 5)- [5(y-1) - 3y + 3]

Answer by stanbon(75887)

25 - [2 + 5y - 3(y + 2)]= -3(2y - 5)- [5(y-1) - 3y + 3]
Work from the inside to the outside.
25 - [2 + 5y - 3y - 6]= -6y +15- [5y-5 - 3y + 3]
25 - [2y-4]= -6y +15- [2y-2]
25 -2y+4 = -6y +15-2y+2
-2y+29 = -8y+17
6y = -12
y = -2
=================
Cheers,
Stan H.